OpenStudy (anonymous):

Which equation represents a function? y2 + x = 6 y4 + x3 = 3 y – x2 = 4 y2 – x2 = 7 y2 + x2 = 7

OpenStudy (anonymous):

I believe the fisrt one but I dojnt know if im right

OpenStudy (welshfella):

no y - x2 = 4 is not a function because some values of map to more that 1 value of x for example y = x2 + 4 when y = 8 x can be 2 or -2

OpenStudy (welshfella):

Yes its the first one because for each value of y there is only one value of x